# Nonlinear Fore(Back)casting and Innovation Filtering for Causal-Noncausal VAR Models

### v3

@@ 7 Concluding Remarks
$\frac{\partial b(v)}{\partial v^{\prime}}=\left(\begin{array}[]{cc}\frac{\partial b_{1}(v)}{\partial v_{1}}&\frac{\partial a_{1}(v)}{\partial v_{2}}\\ \frac{\partial b_{2}(v)}{\partial v_{1}}&\frac{\partial b_{2}(v)}{\partial v_{2}}\end{array}\right)\equiv\left(\begin{array}[]{cc}\cos\theta(v)&-\sin\theta(v)\\ \sin\theta(v)&\cos\theta(v)\end{array}\right).$
Thus the standard identification issue known in the linear SVAR model that is up to a given rotation matrix is replaced by the analogue in which the rotation matrix is local and depends on $v$.
We deduce that:
$\frac{\partial b_{1}(v)}{\partial v_{1}}$ $=$ $\frac{\partial b_{2}(v)}{\partial v_{2}},$
$\frac{\partial b_{1}(v)}{\partial v_{2}}$ $=$ $-\frac{\partial b_{2}(v)}{\partial v_{1}}.$ (a.9)
Let us differentiate the first equation with respect to $v_{1}$ and the second one with respect to $v_{2}$. We get:
$\frac{\partial^{2}b_{1}(v)}{\partial v_{1}^{2}}=\frac{\partial^{2}b_{2}(v)}{\partial v_{1}\partial v_{2}}\;\;\mbox{and}\;\;\frac{\partial^{2}b_{1}(v)}{\partial v_{2}^{2}}=-\frac{\partial^{2}b_{2}(v)}{\partial v_{1}\partial v_{2}},$ (a.10)
and by adding these equalities:
$\frac{\partial^{2}b_{1}(v)}{\partial v_{1}^{2}}+\frac{\partial^{2}b_{1}(v)}{\partial v_{2}^{2}}=0.$ (a.11)
Therefore $b_{1}$ is a harmonic function that satisfies the Laplace equation (a.10). Similarly, $b_{2}$ is also a harmonic function. QED
Harmonic functions are regular functions: they are infinitely differentiable and have series representations that can be differentiated term by term [Axler et al. (2001)]:
$b_{1}(v)$ $=$ $\sum_{h=0}^{\infty}\sum_{k=0}^{\infty}(b_{1hk}v_{1}^{k}v_{2}^{h}),$
$b_{2}(v)$ $=$ $\sum_{h=0}^{\infty}\sum_{k=0}^{\infty}(b_{2hk}v_{1}^{k}v_{2}^{h}).$ (a.12)
Moreover, these series representations are unique. Then, we can apply the conditions (a.8) to these expansions to derive the constraints on the series coefficients and the link between functions $b_{1}$ and $b_{2}$.
Let us define:
$\frac{\partial b_{1}(v)}{\partial v_{1}}=\frac{\partial b_{2}(v)}{\partial v_{2}}\equiv\sum_{h=0}^{\infty}\sum_{k=0}^{\infty}(c_{hk}v_{1}^{h}v_{2}^{k}).$
Then, by integration, we get:
$b_{1}(v)$$\equiv$ $\sum_{h=0}^{\infty}\sum_{k=0}^{\infty}[c_{hk}\frac{v_{1}^{h+1}}{h+1}v_{2}^{k}]+\sum_{k=0}^{\infty}d_{1k}v_{2}^{k},$
$b_{2}(v)$$\equiv$ $\sum_{h=0}^{\infty}\sum_{k=0}^{\infty}[c_{hk}v_{1}^{h}\frac{v_{2}^{k+1}}{k+1}]+\sum_{h=0}^{\infty}d_{2h}v_{1}^{h},$
where the second sums on the right hand sides are the integration ”constants”. Equivalently, we have:
$b_{1}(v)$$\equiv$ $\sum_{k=0}^{\infty}d_{1k}v_{2}^{k}+\sum_{h=1}^{\infty}\sum_{k=0}^{\infty}[c_{h-1,k}\frac{v_{1}^{h}}{h}v_{2}^{k}],$
$b_{2}(v)$$\equiv$ $\sum_{h=0}^{\infty}d_{2h}v_{1}^{h}+\sum_{h=0}^{\infty}\sum_{k=1}^{\infty}[c_{h,k-1}v_{1}^{h}\frac{v_{2}^{k}}{k}].$
Let us now write the second equality in (a.8), i.e.
$\frac{\partial b_{1}(v)}{\partial v_{2}}=-\frac{\partial b_{2}(v)}{\partial v_{1}}.$
This yields:
$\frac{k+1}{h}c_{h-1,k+1}$ $=$ $-\frac{h+1}{k}c_{h+1,k-1},\;\;h\geq 1,k\geq 1,$ (a.13)
$\frac{1}{h}c_{h-1,1}$ $=$ $-(h+1)d_{2,h+1}\;\;h\geq 1,$
$\frac{1}{h}c_{1,k-1}$ $=$ $-(k+1)d_{1,k+1}\;\;k\geq 1,$
$d_{11}$ $=$ $-d_{21}.$
The set of restrictions (a.12) provides information on the dimension of underidentification. As the dimension concerns functional spaces, we describe it from the series expansions (a.11) and the number of independent parameters $b_{1,h,k},b_{2,h,k}$ with $h+k\leq m$. This number is equal to $(m+1)(m+2)/2$.
Proposition A.2
The space of parameters $(b_{1,h,k},b_{2,h,k},\;h+k\leq m)$ is of dimension $2m$.
Proof:
Let us consider an alternative parametrization of (a.12) with parameters $c_{h,k},d_{1,h},d_{2,h}$. The parameters $b_{1,h,k},b_{2,h,k}$ with $h+k=j$ are linear functions of parameters $c_{h,k},h+k=j+1,d_{1,j+1}d_{2,j+1}$. Then the result result follows from restrictions (a.12).
QED
Other identification issues can arise if transformation $b$ is not assumed twice continuously differentiable. Let us consider the first component $v_{1}$ that follows the uniform distribution on $[0,1]$ and introduce two intervals $[0,c]$ and $[1-c,1]$ with $c<0.5$. Then, the variable $w_{1}$ defined by:
$w_{1}=\left\{\begin{array}[]{l}v_{1},\;\mbox{if}\;v_{1}\in(c,1-c),\\ 2v_{1}-1,\;\mbox{if}\;v_{1}\in(0,c)\cup(1-c,1),\end{array}\right.$
also follows the uniform distribution and, similarly to $v_{1}$, variable $w_{1}$ is independent of $v_{2}=w_{2}$.
Note that this transformation is not monotonous. Therefore, the size $\delta$ of a shock to $v_{1}$ is difficult to interpret in terms of a magnitude of a shock to $w_{1}$.
We conclude that, in a nonlinear dynamic framework, the assumption of independence between the components of $v_{t}$ is insufficient to identify the structural innovations to be shocked.

### v4

@@ 7 Concluding Remarks
$\hat{l}_{T}(y|\underline{Y}_{T})=\frac{\hat{l}_{2,T}\left[\hat{A}^{2}\left(\begin{array}[]{c}y\\ \tilde{Y}_{T}\end{array}\right)\right]}{\hat{l}_{2,T}\left[\hat{A}^{2}\left(\begin{array}[]{c}Y_{T}\\ \tilde{Y}_{T-1}\end{array}\right)\right]}\,|\det\hat{J}_{2}|\,\hat{g}_{T}(y-\hat{\Phi}_{1}Y_{T}-\cdots-\hat{\Phi}_{p}Y_{T-p+1}),$
This formula is easily extended to bandwidths adjusted for each component by replacing for example $\frac{1}{h_{1}^{m}}K_{m}\left(\frac{\varepsilon-\hat{\varepsilon}_{t}}{h_{1}}\right)$ by $\prod_{j=1}^{m}\frac{1}{h_{1j}}K\left(\frac{\varepsilon_{j}-\hat{\varepsilon}_{j,t}}{h_{1j}}\right)$, where $K$ is a univariate kernel. Such an adjustment can account for different component variances.
Let us consider the example of a bivariate VAR(1) process with one noncausal component and a scalar noncausal eigenvalue $\lambda_{2}$ (see Section 6). The estimated coefficients of the inverse of $A$ are denoted by:
$\hat{A}^{-1}=\left(\begin{array}[]{cc}\hat{a}^{11}&\hat{a}^{12}\\ \hat{a}^{21}&\hat{a}^{22}\end{array}\right).$
The predictive density depends on unknown scalar parameters $\lambda_{1},\lambda_{2}$ and functional parameters $l_{2},g$ that can be estimated. The marginal density $l_{2}(A^{2}y)$ can be approximated by a kernel estimator:
$\hat{l}_{2,T}(\hat{A}^{2}y)=\frac{1}{T}\frac{1}{h_{2}}\sum_{t=1}^{T}K\left(\frac{\hat{a}^{21}(y_{1}-y_{1,t})+\hat{a}^{22}(y_{2}-y_{2,t})}{h_{2}}\right),$
while the density $l_{2}(A^{2}Y_{T})$ can be approximated by a kernel estimator:
$\hat{l}_{2,T}(y_{T})=\frac{1}{T}\frac{1}{h_{2}}\sum_{t=1}^{T}K\left(\frac{\hat{a}^{21}(y_{1,T}-y_{1,t})+\hat{a}^{22}(y_{2,T}-y_{2,t})}{h_{2}}\right),$
where $h_{2}$ is a bandwidth. The joint density $g(y-\Phi y_{T})$ can be approximated by
$\hat{g}_{T}(y-\hat{\Phi}y_{T})=\frac{1}{T}\frac{1}{h_{11}h_{12}}\sum_{t=1}^{T}K\left(\frac{y_{1}-\hat{\phi}_{1,1}y_{1,T}-\hat{\phi}_{1,2}y_{2,T}-\hat{\varepsilon}_{1,t}}{h_{11}}\right)K\left(\frac{y_{2}-\phi_{2,1}y_{1,T}-\phi_{2,2}y_{2,T}-\hat{\varepsilon}_{2,t}}{h_{12}}\right).$
where $\hat{\varepsilon}_{1,t}$ and $\hat{\varepsilon}_{2,t}$ are residuals $\hat{\varepsilon}_{t}=y_{t}-\hat{\Phi}y_{t-1}$ and $h_{11},h_{12}$ are two bandwidths adjusted for the variations of $\hat{\varepsilon}_{1,t}$ and $\hat{\varepsilon}_{2,t}$, respectively. We get:
$\hat{l}_{T}(y_{1},y_{2}|Y_{T})$ $=$ $\frac{\frac{1}{T}\frac{1}{h_{2}}\sum_{t=1}^{T}K\left(\frac{\hat{a}^{21}(y_{1}-y_{1,t})+\hat{a}^{22}(y_{2}-y_{2,t})}{h_{2}}\right)}{\frac{1}{T}\frac{1}{h_{2}}\sum_{t=1}^{T}K\left(\frac{\hat{a}^{21}(y_{1,T}-y_{1,t})+\hat{a}^{22}(y_{2,T}-y_{2,t})}{h_{2}}\right)}$
×$\times\;|\hat{\lambda}_{2}|\frac{1}{T}\frac{1}{h_{11}h_{12}}\sum_{t=1}^{T}K\left(\frac{y_{1}-\hat{\phi}_{1,1}y_{1,T}-\hat{\phi}_{1,2}y_{2,T}-\hat{\varepsilon}_{1,t})}{h_{11}}\right)$
×$\times\;K\left(\frac{y_{2}-\hat{\phi}_{2,1}y_{1,T}-\hat{\phi}_{2,2}y_{2,T}-\hat{\varepsilon}_{2,t})}{h_{12}}\right).$
C: Nonparametric Estimation of Conditional Quantiles of Filtered Variables
We derive the closed form expression of the estimator of the conditional quantile based on observations $Z_{2,t},t=1,...,T$. When $Z_{2,t}$ is not observed but filtered, the estimator is modified by replacing in the formula $Z_{2,t}$ by $\hat{Z}_{2,t}$.
i) Conditional cdf
The conditional cdf $F_{2}(x,z)=P[Z_{2,t}\leq x|Z_{2,t-1}=z]$ is estimated by the Nadaraya-Watson estimator:
$\hat{F}_{2}(x,z)=\sum_{t=2}^{T}\left[1_{Z_{2,t}\leq x}K\left(\frac{Z_{2,t}-z}{h}\right)\right]/\sum_{t=2}^{T}K\left(\frac{Z_{2,t}-z}{h}\right),$
where $K(.)$ denoted the kernel and $h$ the bandwidth.
This estimated conditional cdf is a piecewise function, with values obtained as follows:
step 1: rank the observations $Z_{2,t},t=2,...,T$, as $Z_{2,(h)},h=1,...,T-1$  [^13] We assume that these values are distinct. Otherwise, we make them distinct by adding a small positive constant..
step 2: The values of $\hat{F}_{2}(x,z)$ are $p(h,z)=\hat{F}_{2}(Z_{2,(h)},z),h=1,...,T-1$. This function is strictly increasing in $h$.
ii) Conditional Quantile
We invert the conditional cdf and evaluate its inverse at probability $\alpha=\Phi(v_{2})$. We get the following closed form estimator
$\hat{G}_{2}(z,v_{2})=Z_{2,k(z,v_{2})}$ where:
$k(z,v_{2})=\left\{\begin{array}[]{ll}1,&\mbox{if}\;\;\Phi(v_{2})<p(1,z).\\ inf\{h:p(h,z)<\Phi(v_{2})\},&\mbox{if}\;\;\Phi(v_{2})>p(1,z).\end{array}\right.$
D: Additional Simulation Results
This Appendix presents additional results on predictions and prediction intervals based on the simulated bivariate process displayed and estimated in Section 6.1. We consider an extend simulated path of 600 observations displayed in Figure a.1:
Figure a.1: Bivariate Mixed VAR(1) Process: $Y_{1}$: solid line, $Y_{2}$: dashed line
Next, we compute oos a series of one-step ahead forecasts starting from $Y_{500}$, based on the estimators obtained from the first 500 observations. The series of 100 one-step ahead forecasts along the trajectory is displayed in Figure a.2.
(a) Estimated 1-step-ahead Forecasts of $Y_{1,t}$
(b) Estimated 1-step-ahead Forecasts of $Y_{2,t}$
Figure a.2: Estimation-Based oos One-Step Ahead Forecasts true:solid line, forecast:dashed line
We observe that the predictions are reliable even during the bubble episodes and on spikes. The predictions are also relatively better for series $Y_{1}$ with longer episodes of local trends.
Unconditional Coverage of the Estimated Prediction Interval
Let us now consider two data generating (DGP) processes. The first one is the bivariate mixed VAR(1) illustrated in 6.1 with the autoregressive matrix: $\Phi=\left(\begin{array}[]{cc}0.7&-1.3\\ 0.0&2.0\end{array}\right),$ eigenvalues $\lambda_{1}=0.7$ and $\lambda_{2}=2.0$. In this experiment, the errors have t-student distributions with 3, 6 and 9 degrees of freedom, respectively. The second process is a bivariate VAR(1) with the autoregressive matrix $\Phi=\left(\begin{array}[]{cc}0.9&-0.3\\ 0.0&1.2\end{array}\right),$ and eigenvalues $\lambda_{1}=0.9,\lambda_{2}=1.2$, which are closer to the unit circle. The errors $\varepsilon_{t}$ have t-student distributions with 3, 6 and 9 degrees of freedom.
The simulated bivariate series are of length 100, 500 and 1000 and each DGP is replicated 500 times. The last observation from each simulated path is set aside. It is forecast and used for forecast coverage assessment. The autoregressive parameters $\hat{\Phi}$ are estimated by the GCov estimator from the autocovariances of errors $\varepsilon_{t}=Y_{t}-\Phi Y_{t-1}$, and of their second, third and fourth powers, up to lag $H=10$. The algorithm maximizing the GCov objective function is initiated each time at the starting values 0.1, 0.1, 0.5, 0.5. The oos predictive density of $y(T+1)$ given in (3.2) is evaluated from the parameter estimates, over a grid of 100 possible future values for each of the components series. We use Gaussian kernels and bandwidths $h_{2}=s.d.(Y_{2,t})$ $h_{11}=s.d.(\varepsilon_{1})$ and $h_{12}=s.d.(\varepsilon_{2})$ (see Online Appendix B for kernel density estimators). The main mode of the predictive density, estimated from formula (3.2) with Gaussian kernels, provides the point forecast. The oos prediction intervals at horizon 1 are obtained from the marginal 10th and 90th percentiles of the bivariate predictive density. As before, the rationale for choosing level 80% is to ensure a sufficiently large number of observations in the tails for computing the quantiles of predictive densities.
The unconditional coverage of the estimated prediction interval is reported in Table a.1 below for the two DGPs described above, sample sizes T=500 and T=1000 and t-student distributions of error $\varepsilon_{t}$ with 3, 6 and 9 degrees of freedom. We observe that it is either
T=100 T=500 T=1000
VAR(1) with eigenvalues 0.7, 2.0
component t(3) t(6) t(9) t(3) t(6) t(9) t(3) t(6) t(9)
$y_{1}(T+1)$ 90.0 91.0 92.6 91.4 92.8 92.0 91.4 90.8 93.2
$y_{2}(T+1)$ 92.2 93.8 91.8 93.4 94.0 94.8 95.2 93.2 94.5
VAR(1) with eigenvalues 0.9, 1.2
component t(3) t(6) t(9) t(3) t(6) t(9) t(3) t(6) t(9)
$y_{1}(T+1)$ 80.2 82.8 87.4 82.2 86.0 84.2 84.6 84.4 88.6
$y_{2}(T+1)$ 90.4 89.4 90.8 90.6 91.6 92.40 91.4 91.0 92.8
Table a.1: Unconditional Coverage of GCov Estimated Prediction Interval at 80%
Based on the sample quantiles of estimated predictive densities from sample sizes of 100, 500 and 1000 observations (col.2:4) for t-student distributed errors with 3,6 and 9 degrees of freedom
greater or close to the theoretical conditional coverage of the prediction interval for both DGPs and sample sizes. It does not mean that the prediction interval is ”conservative”. The reported coverage rate depends not only on the prediction quality but also on the accuracy of the quantile estimator. Because the predictive density is approximated at a relatively low number of points (10000), has long tails and the tail observations are not dense, we use a robust quantile estimator, which approximates the quantiles ”from above”, resulting in practice in a unconditional coverage higher than the nominal conditional coverage rate.
Estimated Prediction Set Uncertainty
This Section illustrates the uncertainty in the estimated conditional prediction interval described in Section 4. We consider the simulated trajectory of length T=200 of the bivariate VAR(1) process illustrated in Table 1 with autoregressive matrix: $\Phi=\left(\begin{array}[]{cc}0.9&-0.3\\ 0.0&1.2\end{array}\right),$ and eigenvalues $\lambda_{1}=0.9,\lambda_{2}=1.2$ and t(6) distributed errors with an identity variance-covariance matrix. We are interested in the conditional prediction interval out-of-sample of the first component $Y_{1,T+1}$ = $Y_{1,200}$ given the past and current values of the two series.
The mixed VAR(1) model is estimated by the GCov estimator from the observations $t=1,...,199$ with four power transforms of model errors and lag $H=10$, providing the following estimates of autoregressive parameters: 0.8931, -0.2146, 0.0180, 1.2797, and eigenvalues: 0.903 and 1.269. The estimated predictive density of $Y_{200}$ conditional on $Y_{199}$ is computed from formula (3.2) by using kernel smoothed density estimators. More specifically, we employ Gaussian kernels and bandwidths $h_{2}=s.d.(Y_{2}),h_{11}=s.d.(\varepsilon_{1}),h_{12}=s.d(\varepsilon_{2})$ (see Online Appendix B for the kernel density estimators). For $\alpha_{1}=0.2$ and $\Phi^{-1}(\alpha_{1}/2)=-1.28$, the estimated prediction interval of $Y_{1,T+1}=Y_{1,200}$ is: $\widehat{PI}(y,\alpha_{1})=[-6.954,0.360]$. It contains the true value of $Y_{1,200}=-0.696$.
In the next step, we replicate the initial paths $S=50$ times by backcasting from the bivariate terminal condition $Y_{199}=[-1.188,0.473]$. We use the backcasting formula given in Corollary 2, evaluated from estimated model parameters and kernel density estimators. We employ Gausian kernels and bandwidths $h_{1}=s.d.(Y_{1}),h_{11}=s.d.(\varepsilon_{1})$ and $h_{12}=s.d.(\varepsilon_{2})$.
The randomness of the backcast paths conditional on $Y_{T}$ is generated as follows: i) We first draw 100 values as if the components $Y_{1,T-1},Y_{2,T-1}$ were independent conditional on $Y_{T}$. This is done by inverting the estimated conditional c.d.f.s of $Y_{1,T-1}$ and $Y_{2,T-1}$ given $Y_{T}$. This provides the sampling with the importance (misspecified) density; ii) Next, we re-sample in the set of values obtained in step i) above, with the weights proportional to the ratio of the joint backward predictive density divided by the product of the two marginal backward predictive densities. This procedure adjusts for the omitted cross-sectional dependence in step i).
The parameter $\Phi$ and functional parameter $g$ are re-estimated from each replicated path. Next, the values of $\hat{m}^{s}$, $\hat{\sigma}^{s}$ are computed from the quantiles of $50$ predictive densities of $Y_{T+1}$ conditional on $Y_{T}$. Then, we obtain the bootstrap confidence interval at level $0.951-\alpha_{2}=$ 0.95 of the prediction interval $\widehat{CSPI}(y_{1},\alpha_{1},\alpha_{2})=[-11.839,5.246],$ with the solution $\hat{q}(y,\alpha_{1},\alpha_{2})=2.99$. The confidence interval $\widehat{CSPI}$ of the prediction interval is much larger than the interval $\widehat{PI}$. This result illustrates the possibility of extending the level of the prediction interval from 80% to 95% while taking into account the estimation risk on $\Phi$ and $g$ given that the estimator of the functional parameter $g$ converges at a lower speed than the matrix parameter estimator. The extended interval is such that the length of the interval has almost doubled and the interval became less symmetric with respect to 0.
E: Additional Empirical Results
1. Oil Prices and US GDP Rates Estimation
The quarterly macroeconomic time series are displayed in Figure a.3.
Figure a.3: The path of Oil and US GDP rate with prediction dates
The green vertical lines indicate the prediction points discussed in the main text.
1.1 Mixed VAR(1): Estimation results
The estimated errors of the mixed VAR(1) model are plotted in Figure a.4.
Figure a.4: Errors $\hat{\epsilon}_{t}$, Noncausal VAR(1), Q1 1986 -Q2 2019
We observe that their densities are non-Gaussian, as shown in Figure a.5:
(a) Density of $\hat{\epsilon}_{1,t}$
(b) Density of $\hat{\epsilon}_{2,t}$
Figure a.5: Density of $\hat{\epsilon}_{t}$
(a) ACF of $\hat{\varepsilon}$
(b) ACF of $\hat{\varepsilon}^{2}$
(c) ACF of $\hat{\varepsilon}^{3}$
Figure a.6: Residual ACF
We find that the estimated residuals, their squares and third powers are serially uncorrelated (Figure a.6). The error variance-covariance matrix is $\hat{\Sigma}=\left[\begin{array}[]{cc}0.262&0.184\\ 0.184&1.506\end{array}\right]$ and contemporaneous correlation is 0.28 (statistically significant at 0.05).
2. Predictive Densities
We provide the plots of predictive densities of the estimated mixed VAR(1) model of oil prices and US GDP rate computed oos, one step-ahead and conditional on selected dates between T=112 and T=118 marked by the shaded area in Figures a.7-a.11.
Figure a.7: VAR(1) based predictive density estimation, T=112
Figure a.8: VAR(1) based predictive density estimation, T=114
Figure a.9: VAR(1) based predictive density estimation, T=115
Figure a.10: VAR(1) based predictive density estimation, T=116
Figure a.11: VAR(1) based predictive density estimation, T=118
We observe deformations during the bubble episode arising due to increasing downturn probabilities, which is consistent with the patterns of univariate predictive densities documented in Hencic et al. (2020), Hecq et al. (2022) and the theoretical results of Freis, Zakoian (2019).